# An Empirical Analysis of Compute-Optimal Inference for Problem-Solving with Language Models

### v1

@@ 3.2.3 Reward Balanced Search (REBASE)
The REBASE tree search method, illustrated in Fig. 3, inherits the exploitation and pruning properties of tree search, while using the reward model alone to estimate the nodes’ qualities without additional computation for estimating values by sampling children. The efficiency is achieved by constraining the total expansion width of the tree at a certain depth. REBASE balances the expansion width among the nodes at the same depth based on the rewards given by the Process Reward Model (PRM). The details are provided below:
Notations. We view the fine-tuned LLM as a policy $\pi_{\theta}$ which generates the solution step by step. Given a question $x$ and the first $k$ steps of a solution $r_{1}\cdots r_{k}$, the $(k+1)$-th step is sampled from $\pi_{\theta}(\cdot|xr_{1}\cdots r_{k})$. REBASE effectively generates a solution tree during inference, in which the root node the question $x$ and other nodes corresponds to solution steps. When generating solution trees, we generate children of $r_{k}$ by sampling from $\pi_{\theta}(\cdot|xr_{1}\cdots r_{k})$. Here we slightly abuse notations and use the corresponding question/solution step to denote a node. The reward of a node $r_{k}$ is generated by the PRM: $R(r_{k}):=R(qr_{1}\cdots r_{k})$.
Initialization. Given the question $x$, balance temperature $T_{b}>0$, and sampling number of solutions $N$, we sample $N$ instances of the first step for the question, yielding all the nodes of depth 1 in the search tree. We let the sampling budget of depth 0, $B_{0}$, to $N$ at initialization.
Reward modeling and update. In the $i$-th iteration, the PRM assigns the rewards to all the nodes at depth $i$. After that, the algorithm examines whether the solutions up to depth $i$ are complete. Supposing there are $C_{i}$ completed solutions, we update the sampling budget using $B_{i}\leftarrow B_{i-1}-C_{i}$. If $B_{i}=0$, the process ends, and we obtain $N$ solutions.
@@ 4 Experiments
Figure 4: The inference computation scaling comparisons across model sizes. The left/right panel shows the problem-solving error rate on MATH based on Weighted Mjority/Best-of-N.
Figure 5: The inference computation scaling comparisons across model sizes. The left/right panel shows the problem-solving error rate on GSM8K based on Weighted Mjority/Best-of-N. MCTS is not included in the comparison because of its poor compute-accuracy trade-off.
@@ 4.1 Setup
We conduct experiments on two mathematical problem-solving datasets to investigate the scaling effects of computation and our REBASE method for both challenging and simpler problems. Specifically, MATH (Hendrycks et al., 2021a) and GSM8K (Cobbe et al., 2021b) are datasets containing high school mathematics competition-level problems and grade-school level mathematical reasoning problems, respectively. Following (Lightman et al., 2023b; Wang et al., 2024; Sun et al., 2024), we use the MATH500 subset as our test set.
Policy model (generators).
To study the inference compute scaling effect, we choose Pythia (Biderman et al., 2023) as our base models since various model sizes are available in the Pythia family. For tree search, we use math-specialized Llemma models (Azerbayev et al., 2024). We further finetune these models on the MeatMath dataset (Yu et al., 2024) using full parameter supervised fine-tuning (Full-SFT), The detailed finetuning configuration is given in the Appendix. Additionally, we test the Mistral-7B (Jiang et al., 2023) to expand our findings across different models and architectures.
@@
We use sampling and tree search methods to generate multiple samples and select the answer through Best-of-N, Majority Voting, or Weighted Voting. Each configuration has been run multiple times to calculate the mean and variance, thereby mitigating the randomness and ensuring the reliability of our conclusions.
4.2 Main Results of Compute-Optimal Inference
@@
In order to compare the compute budgets of different models, we plot the figures with the number of FLOPs used per question during inference. We compute the inference FLOPs based on the standard formula from (Kaplan et al., 2020).
Scaling law of compute-optimal inference.
@@
The scaling effect of inference computation budget across different model sizes is shown in Fig. 1. We note that the error rate first decreases steadily and then starts to saturate. Where smaller model first takes advantage since it can generate more samples given limited budget, larger models are preferable with more FLOPs due to the saturation of small model performances. As highlighted in the right panel of Fig. 1, the optimal model size can be different under various computation budgets. We performed a regression analysis on inference FLOPs $C$ and corresponding model sizes $N$ to establish a relationship between a given computational budget and its optimal model size. The resulting regression equation, $\log_{10}\left(C\right)=1.19\log_{10}\left(N\right)+2.03$, enables us to estimate the optimal inference model size for a specified computational constraint.
Llemma-7B model achieves competitive accuracy to Llemma-34B model with lower compute budget.
Fig. 4 & 5 show the curves of error rates versus total number of inference FLOPs per question. Inference methods with different model sizes are plotted in the same diagram. We found that Llemma-7B costs approximately 2×2$\times$2 × less total FLOPs than Llemma-34B under the same method (Sampling, MCTS, REBASE) and task (MATH, GSM8K) while achieving competitive accuracy. This result suggests that, with the same training dataset and model family, training and inference with a smaller model could be more favorable in terms of compute budget if multiple sampling or search methods are employed.
@@
Figure 6: The inference computation scaling laws of different models for the problem-solving error rate on MATH500 test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, W.M. and BoN refer to Weighted Majority and Best-of-N, respectively.
@@
Figure 7: The inference computation scaling laws of different models for the problem-solving error rate on GSM8K test set. The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, W.M. and BoN refer to Weighted Majority and Best-of-N, respectively.
@@
Table 1: REBASE with lower compute budget has competitive accuracy against Sampling with higher compute budget. We use weighted voting to aggreagte the candidate solutions in both Sampling and REBASE.
@@
4.3 Comparing REBASE to Other Baselines
REBASE is Pareto-optimal.
While MCTS undeperforms Sampling (Fig. 4), from Fig. 4, 5, 6, and 7, we found that REBASE consistently outperforms the Sampling method in all settings, when fixing the model and the evaluation task. Table 1 shows that REBASE can achieve competitive accuracy with even a lower compute budget than the sampling method. This finding is novel, and differs from previous tree search works which typically improve the performance at the cost of higher computational expense compared to sampling (Chen et al., 2024a; Xie et al., 2023). Table 2 shows that given the same compute budget (sampling 32 solutions for the 7B model and 8 solutions for 34B model), using REBASE yields higher accuray than sampling.
Table 2: Accuracy of diffrent inference configurations under a specific compute budget. MV, BoN and WV denote Majority Voting, Best-of-N and Weighted Voting, respectively.
# Samples MATH FLOPs GSM8K FLOPs MATH500 GSM8K
Mistral-7B
Greedy 1 $3.4\times 10^{12}$ $2.3\times 10^{12}$ 28.6 77.9
Sampling + MV 32 $109.2\times 10^{12}$ $72.6\times 10^{12}$ 36.1 85.7
Sampling + BoN 32 $109.2\times 10^{12}$ $72.6\times 10^{12}$ 40.3 89.4
Sampling + WV 32 $109.2\times 10^{12}$ $72.6\times 10^{12}$ 39.7 89.1
REBASE + MV 32 $136.2\times 10^{12}$ $78.9\times 10^{12}$ 44.1 88.8
REBASE + BoN 32 $136.2\times 10^{12}$ $78.9\times 10^{12}$ 45.4 89.4
REBASE + WV 32 $136.2\times 10^{12}$ $78.9\times 10^{12}$ 45.0 89.8
Llemma-7B
Greedy 1 $3.92\times 10^{12}$ $2.3\times 10^{12}$ 30.0 68.5
Sampling + MV 32 $125.4\times 10^{12}$ $73.9\times 10^{12}$ 41.0 80.0
Sampling + BoN 32 $125.4\times 10^{12}$ $73.9\times 10^{12}$ 41.7 85.6
Sampling + WV 32 $125.4\times 10^{12}$ $73.9\times 10^{12}$ 43.5 85.4
REBASE + MV 32 $148.0\times 10^{12}$ $82.6\times 10^{12}$ 46.1 86.1
REBASE + BoN 32 $148.0\times 10^{12}$ $82.6\times 10^{12}$ 44.1 86.9
REBASE + WV 32 $148.0\times 10^{12}$ $82.6\times 10^{12}$ 46.8 87.3
Llemma-34B
Greedy 1 $19.0\times 10^{12}$ $11.2\times 10^{12}$ 33.0 78.4
Sampling + MV 8 $152.3\times 10^{12}$ $89.7\times 10^{12}$ 39.9 84.3
Sampling + BoN 8 $152.3\times 10^{12}$ $89.7\times 10^{12}$ 40.4 86.7
Sampling + WV 8 $152.3\times 10^{12}$ $89.7\times 10^{12}$ 41.0 86.0
REBASE + MV 8 $176.8\times 10^{12}$ $98.7\times 10^{12}$ 43.9 86.1
REBASE + BoN 8 $176.8\times 10^{12}$ $98.7\times 10^{12}$ 43.6 86.9
REBASE + WV 8 $176.8\times 10^{12}$ $98.7\times 10^{12}$ 42.9 86.9
Weaker models gain more from Tree Search.
For example, our proposed REBASE leads to $5.3\%$, $3.3\%$, and $2.6\%$ performance gains on MATH for Mistral-7B, Llemma-7B, Llemma-34B, respectively. The order of accuracy increase is inversely related to the model’s corresponding greedy search on those datasets. This suggests that weaker models, as indicated by their lower greedy search accuracy, benefit more from tree search methods like REBASE.
REBASE saturates later than sampling with higher accuray.
From Fig. 6 and Fig. 7, we observe that both sampling and REBASE saturate early in GSM8K and relatively late in MATH, which we attribute to the difference of the difficulty level. This can be explained through the LLM may assign high probability to the true answer in easy problems than those of harder problems, as suggested by Theorems 1 & 2 with their proofs A. On MATH (Fig. 6), we see that REBASE finally saturates with a higher accuracy than sampling. We hypothesize the reason is that REBASE samples the truth answer with higher probability than sampling. And as demonstrated by Theorems 1 & 2, the upper bound becomes higher.
5 Conclusions & Limitations
@@ Conclusions.
In this work, we conducted a comprehensive empirical analysis of compute-optimal inference for problem-solving with language models. We examined the scaling effect of computation during inference across different model sizes and found that while increased computation generally leads to higher performance, the optimal model size varies with the available compute budget. When the computation budget is limited, smaller models can be preferable. Specifically, a small language model can achieve competitive accuracy to a model four times larger while using approximately half the total FLOPs under the same inference methods (Sampling, MCTS, REBASE) and tasks (MATH, GSM8K). Additionally, we introduce our novel tree search method, REBASE, which is more compute-optimal than both sampling methods and Monte Carlo Tree Search (MCTS). REBASE typically achieves higher accuracy while using several times less computation than sampling methods. Our results underscore the potential of deploying smaller models equipped with sophisticated inference strategies like REBASE to enhance problem-solving accuracy while maintaining computational efficiency.
@@
Our experiments specifically targeted mathematical problem-solving tasks. Additionally, our models were primarily trained using the MetaMath dataset. Investigating the impact of different training datasets on the performance and efficiency of compute-optimal inference strategies for mathematical problem-solving would be a valuable direction for future research.
Acknowledgement.
The computation of the experiments is partially supported by Netmind.
@@ A.1 Proof of Theorem 1
${\mathbb{P}}(E_{n})\leq$ ${\mathbb{P}}(\exists\leavevmode\nobreak\ y^{\prime\prime}\in{\mathcal{A}}\backslash\{y^{*}\},f_{n}(y^{\prime\prime})\geq f_{n}(y^{*}))$
@@
$\exists\leavevmode\nobreak\ N\in{\mathbb{N}}^{*},\text{ such that for any }n\geq N,\leavevmode\nobreak\ y^{*}=\operatorname*{arg\,max}_{y\in{\mathcal{A}}}f_{n}(y)$
@@ Appendix C Hyper-parameters
Appendix D Supplementary Figures
@@
We append the figures about Majority Voting and Majority Voting v.s. Weighted Majority Voting (Fig. 8, 9 ,10, 11) in this section. The experiments show that although the gap between Majority Voting and Weighted Majority Voting on sampling is huge. This gap becomes much smaller if we apply REBASE. This phenomenon can be caused by the selection ability of tree search like REBASE. Once REBASE already samples solutions with high rewards, conducing Weighted Majority Voting gains less since the sampled solutions may all have relatively high and stable rewards compared with those of sampling.

### v2

@@ 3.1.2 Reward Balanced Search (Rebase)
The Rebase tree search method, illustrated in Fig. 3, inherits the exploitation and pruning properties of tree search, while using a reward model alone to estimate quality of intermediate nodes. This saves on computation compared to methods such as MCTS, since it does not involve estimate node quality with explicit rollouts. In short, the underlying idea is to use a process reward model to determine how much each node should be expanded at each depth. Namely, Rebase expands nodes at a given depth according to their softmax-normalized reward scores, subject to a total expansion budget. We describe this procedure in more detail below.
Notations. We view the fine-tuned LLM as a policy $\pi_{\theta}$ which generates the solution step by step. Given a question $x$ and the first $k$ steps of a solution $r_{1}\cdots r_{k}$, the $(k+1)$-th step is sampled from $\pi_{\theta}(\cdot|xr_{1}\cdots r_{k})$. Rebase generates a solution tree during inference, in which the root node is the question $x$, and other nodes corresponds to solution steps. When generating solution trees, we generate children of $r_{k}$ by sampling from $\pi_{\theta}(\cdot|xr_{1}\cdots r_{k})$. We use the corresponding solution step to denote a node. The reward of a node $r_{k}$ is generated by the PRM: $R(r_{k}):=R(qr_{1}\cdots r_{k})$.
Initialization. Given the question $x$, a balance temperature $T_{b}>0$, and target number of generated solutions $N$, we sample $N$ instances of the first step for the question, yielding all the nodes of depth 1 in the search tree. We let the sampling budget of depth 0, $B_{0}$, to $N$ at initialization.
Reward assignment and update. In the $i$-th iteration, the PRM assigns the rewards to all the nodes at depth $i$. After that, the algorithm examines whether the solutions up to depth $i$ are complete. Supposing there are $C_{i}$ completed solutions, we update the sampling budget using $B_{i}\leftarrow B_{i-1}-C_{i}$. If $B_{i}=0$, the process ends, and we obtain $N$ solutions.
@@ 4 Experiments
Our experiments are centered around two main questions:
  • Compute-optimal model size: How does performance scale as inference-time compute is increased with a fixed inference strategy, but with varying model size?
  • Compute-optimal inference strategy: How does performance scale as inference-time compute is increased with various inference strategies (and various model sizes)?
We detail our experimental setup below.
Figure 4: MATH inference scaling across inference strategies and model sizes (lower is better). The left/right panel shows the error rate on MATH based on weighted majority/best-of-$n$. Rebase is the compute-optimal strategy at all budgets, with 7B typically the optimal model size.
Figure 5: GSM8k inference scaling across inference strategies and model sizes (lower is better). The left/right panel shows the problem-solving error rate on GSM8K based on weighted majority/best-of-$n$. MCTS is not included in the comparison because of its poor compute-accuracy trade-off. Rebase is the compute-optimal inference strategy, and the optimal model size varies.
@@ 4.1 Setup
We conduct experiments on two mathematical problem-solving datasets to investigate the effects of scaling inference compute for both challenging and simpler problems. Specifically, MATH (Hendrycks et al., 2021a) and GSM8K (Cobbe et al., 2021b) are datasets containing high school mathematics competition-level problems and grade-school level mathematical reasoning problems, respectively. Following (Lightman et al., 2023b; Wang et al., 2024; Sun et al., 2024), we use the MATH500 subset as our test set.
Policy model (solution generator).
To study the how performance scales as inference compute is increased using a fixed strategy, the primary axis of variation is model size. Therefore, we choose Pythia (Biderman et al., 2023) as our base models, since various model sizes are available in the Pythia family. To study inference scaling under different inference strategies (e.g., tree search, weighted majority voting), we use math-specialized Llemma models (Azerbayev et al., 2024). We finetune these models on the MetaMath dataset (Yu et al., 2024) using full parameter supervised fine-tuning (Full-SFT), The finetuning configuration is given in the Appendix. Additionally, we test the Mistral-7B (Jiang et al., 2023) to expand our findings across different models and architectures.
@@
We use sampling and tree search methods to generate multiple candidates, and select the answer through best-of-$n$, majority voting, or weighted voting. Each configuration is run multiple times to calculate the mean and variance, which mitigates effects from randomness and thereby improves the reliability of our conclusions.
4.2 Compute-Optimal Model Size
@@
To compare the inference compute budgets of different models, we plot the figures with the number of FLOPs used per question during inference. We compute the inference FLOPs based on the standard formula from (Kaplan et al., 2020).
Scaling law of compute-optimal inference for model size.
@@
Fig. 1 shows the relationship between inference compute and error rate for different model sizes. The error rate first decreases steadily and then starts to saturate. Initially, sampling many times from smaller models is compute-optimal. At larger compute budgets the larger models are preferable, since the performance of small models has saturated. As highlighted in the right panel of Fig. 1, the optimal model size varies based on the inference budget. We performed a regression analysis on inference FLOPs $C$ and model sizes $N$ to establish a relationship between a given computational budget and its optimal model size. The resulting equation, $\log_{10}\left(C\right)=1.19\log_{10}\left(N\right)+2.03$, lets us estimate the optimal inference model size for a specific compute budget.
Llemma-7B achieves competitive accuracy to Llemma-34B with less compute.
Fig. 4 and Fig. 5 shows the relationship between error rate and inference FLOPs for Llemma 7B and Llemma 34B using different inference strategies. Llemma-7B requires around 2×2$\times$2 × less total FLOPs than Llemma-34B to achieve comparable accuracy. This held across inference strategies (sampling strategies, MCTS, Rebase) and tasks (MATH, GSM8K). This result suggests that, with the same training dataset and model family, generating more tokens with a suitable inference strategy using a smaller model can have more favorable cost-performance tradeoffs than using a larger model.
@@
Figure 6: MATH inference scaling across inference strategies and models (lower is better). The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, W.M. and BoN refer to weighted majority and best-of-$n$, respectively.
@@
Figure 7: GSM8K inference scaling across inference strategies and models (lower is better). The tested models are Llemma-7B (left), Llemma-34B (middle), & Mistral-7B (right). In the legend, W.M. and BoN refer to weighted majority and best-of-$n$, respectively.
4.3 Compute-Optimal Inference Strategy
Rebase is Pareto-optimal.
Rebase consistently achieves the best cost-performance tradeoffs, outperforming the sampling-based methods in all settings when fixing the model and the evaluation task (Fig. 4, 5, 6, and 7). For example, in Fig. 4, Rebase is the compute-optimal strategy at all inference compute budgets, with 7B typically the optimal model size. On the other hand, MCTS underperforms the sampling-based methods at each compute budget, likely due to its costly rollouts (Fig. 4) compared to the efficient use of the reward model in Rebase.
Table 1 shows that Rebase achieves better accuracy with a lower compute budget compared to sampling-based weighted voting. With the 7B model, Rebase achieves higher accuracy with 7 times less compute. This finding is novel, and differs from previous tree search methods that typically improve the performance at the cost of higher computational expense compared to sampling-based voting (Chen et al., 2024a; Xie et al., 2023).
Weaker models gain more from tree search.
For example, our proposed Rebase leads to $5.3\%$, $3.3\%$, and $2.6\%$ performance gains on MATH for Mistral-7B, Llemma-7B, Llemma-34B, respectively. The order of accuracy increase is inversely related to the model’s corresponding greedy search accuracy on those datasets. This suggests that weaker models, as indicated by their lower greedy search accuracy, benefit more from tree search methods like Rebase.
Rebase saturates later than sampling with higher accuracy.
From Fig. 6 and Fig. 7, we observe that both sampling and Rebase saturate early in GSM8K and relatively late in MATH. We attribute this to the difference of in difficulty levels between GSM8K and MATH. Specifically, the LLM may assign high probability only to correct solutions in easy problems, but spread probability mass across solutions in harder problems. Thus, harder problems may require aggregating over more solution paths to converge to the distribution over answers shown in Theorems 1 & 2. On MATH (Fig. 6), we see that REBASE finally saturates with a higher accuracy than sampling. We hypothesize the reason is that drawing samples from Rebase corresponds to sampling from a policy that assigns high probability to true answers compared to sampling from the underlying language model. If this was indeed the case, Theorems 1 & 2 indicate that the upper bound would become higher. We leave formally analyzing the behavior of tree search algorithms as interesting future work.
@@ 4.3 Compute-Optimal Inference Strategy
Table 1: Rebase with a lower compute budget has better accuracy than sampling with a higher compute budget. We use weighted voting to aggregate candidates for both sampling and Rebase.
@@
5 Conclusions, Limitations, and Discussions on Concurrent Works
@@ Conclusions.
We study the relationship between task performance and the amount of compute expended during inference for various model sizes, model families, and inference strategies, to form empirical inference scaling laws. These relationships let us reason about compute-optimal inference: inference configurations that give the best performance at a given compute budget.
Our results lead to three main takeaways. First, we find that using a smaller model and generating more tokens in an inference strategy often outperforms using a larger model at a fixed compute budget. This has implications for models deployed in the real world, where inference compute is constrained in various ways. Specifically, it is potentially beneficial to deploy smaller models with more sophisticated inference strategies for better cost-performance tradeoffs. Second, we showed that in the limit of infinite compute (allocated by drawing more samples), sampling-based majority voting strategies inevitably saturate to a distribution that depends on the underlying generation policy. Hence, it is of interest to alter the sampling distribution by designing an alternative inference strategy. Third, we design such an inference strategy–the novel Rebase tree search–and find it is Pareto optimal, in that it achieves the best performance across all tested compute budgets. Notably, it outperforms commonly used weighted majority voting and MCTS methods that have attracted much interest and widespread use. This finding not only shows the strength of Rebase, but also indicates that there is large headroom to improve language model performances via inference-time algorithms.
@@
Our empirical analysis specifically targets mathematical problem-solving. Investigating the inference scaling laws and compute-optimal inference strategies for tasks beyond mathematical problem-solving would be a valuable direction for future research. Additionally, we mainly evaluate the proposed Rebase on the GSM8K and MATH500 datasets. We speculate that the Rebase algorithm, which assumes access only to a function that assigns scores to nodes, will be effective in tasks beyond those studied here.
@@ A.1 Proof of Theorem 1
${\mathbb{P}}(E_{n})\leq$ ${\mathbb{P}}(\exists~{}y^{\prime\prime}\in{\mathcal{A}}\backslash\{y^{*}\},f_{n}(y^{\prime\prime})\geq f_{n}(y^{*}))$
@@
$\exists~{}N\in{\mathbb{N}}^{*},\text{ such that for any }n\geq N,~{}y^{*}=\operatorname*{arg\,max}_{y\in{\mathcal{A}}}f_{n}(y)$
@@ Appendix C Hyper-parameters
Appendix D Supplementary Figures and Tables
@@
In this section, we additionally include experimental results on the majority voting method, along with its comparison with weighted majority voting (Fig. 8, 9 ,10, 11). The experiments show that although the gap between majority voting and weighted majority voting on sampling is huge. This gap becomes much smaller if we apply Rebase. This phenomenon can be caused by the selection ability of tree search like Rebase. Once Rebase already samples solutions with high rewards, conducing weighted majority voting gains less since the sampled solutions may all have relatively high and stable rewards compared with those of sampling.